OpenStudy (anonymous):

please help:John is traveling to a meeting that is 28 miles away. He needs to be there in 30 minutes. How fast does he need to go to make it to the meeting on time

OpenStudy (anonymous):

Tham, distance, speed and time follow a basic rule: DISTANCE = SPEED X TIME Hence, with your question: \[28 = speed\times\frac{30}{60}\] \[speed = \frac{28\times60}{30} = 56\text{ }miles\text{ }per\text{ }hour\]